# Robust shallow shadows

### v1

@@ Appendix D Worst-case bias per Pauli support
In Figure 8 we show the same worst-case bias of standard non-mitigated shadow estimation for Pauli observables, under the same noise conditions. This is presented in shaded lines. However, this time we also present the worst-case bias for fixed Pauli supports $\mathrm{PS}=1,2$. We observe that the Pauli worst-case biases for these Pauli supports seem to hit a plateau for both noise levels. However, for $r\sim 10^{-2}$, even though stable, the biases are still sizeable.
@@
Here, we present the distributions of the data used to estimate the eigenvalues of the noisy frame operator obtained on the ibm_nairobi quantum processor using the IBMQ platform. As in Fig. 5 (a), we show one eigenvalue per Pauli support $\operatorname{PS}$ and per depth $D\in\{0,1,2\}$. Moreover, we use a sequence of size |$\absolutevalue{\mathcal{S}}=10^{5}$.
Figure 9: Histograms of experimental realization on IBMQ. Distribution of samples used to estimate eigenvalues of the noisy frame $\mathcal{S}$ for different Pauli supports $1\leq\operatorname{PS}\leq 5$, over $|\mathcal{S}|=10^{5}$ samples from the ibm_nairobi quantum processor, for $n=5$ and $D\in\{0,1,2\}$. The average over the $10^{5}$ samples gives our estimates of the eigenvalues of $\tilde{S}$ shown in Fig. 5 (a). We indicate the corresponding eigenvalue by dashed red lines, while the dashed black lines indicate the $1/(d+1)$ level.

### v2

@@ Appendix D Worst-case bias per Pauli support
In Figure 8 we show the same worst-case bias of standard non-mitigated shadow estimation for Pauli observables, under the same noise conditions. This is presented in shaded lines. However, this time we also present the worst-case bias for fixed Pauli supports $\mathrm{PS}=1,2$. We observe that the Pauli worst-case biases for these Pauli supports seem to hit a plateau for both noise levels. However, for $r\sim 10^{-2}$, even though stable, the biases are still sizable.
@@
Here, we present the distributions of the data used to estimate the eigenvalues of the noisy frame operator obtained on the ibm_nairobi quantum processor using the IBM Quantum platform. As in Fig. 5 (a), we show one eigenvalue per Pauli support $\operatorname{PS}$ and per depth $D\in\{0,1,2\}$. Moreover, we use a sequence of size |$\absolutevalue{\mathcal{S}}=10^{5}$.
We observe that multiple eigenvalues significantly deviate from the value for global rotations still for $D=2$. Furthermore, the bias significantly differs for different eigenvalues. Thus, mitigating the bias by rescaling the estimator with a single “global” scalar or equivalently only relying on ratio estimators does not remove the bias.
Figure 9: Histograms of experimental realization on IBM Quantum. Distribution of samples used to estimate eigenvalues of the noisy frame $\mathcal{S}$ for different Pauli supports $1\leq\operatorname{PS}\leq 5$, over $|\mathcal{S}|=10^{5}$ samples from the ibm_nairobi quantum processor, for $n=5$ and $D\in\{0,1,2\}$. The average over the $10^{5}$ samples gives our estimates of the eigenvalues of $\tilde{S}$ shown in Fig. 5 (a). We indicate the corresponding eigenvalue by dashed red lines, while the dashed black lines indicate the $1/(d+1)$ level.